# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v5

@@ 4.4. GPC Opt-out Evaluation
In addition to our app-level opt-out UI evaluation we performed a platform-level GPC opt-out evaluation. The “selling” or “sharing” of personal information by an app operator corresponds to a third party’s “buying” or “collecting” of personal information. Thus, to determine whether an app is selling or sharing personal information we identified the most prevalent third-party tracking domains across all apps in our app dataset. Our starting point for this task was the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). We found that 33 third-party tracking domains from the X-Ray Tracker List have a 99% coverage with at least one tracking domain in each app in our app dataset. [^20] We combined trackers with different subdomains in our app dataset into one domain to match the tracker domains in the X-Ray Tracker List. For example, we combined app.adjust.com, view.adjust.com, app.us.adjust.com, and s2s.adjust.com into adjust.com. For the 33 third parties associated with these domains we evaluated their privacy policies and developer documentation and looked for explicit references that—per the CCPA—they buy or collect personal information from app operators that integrate their libraries or sell or share such downstream.
@@
To illustrate, Meta’s policy (Meta, 2024) indicates that it is buying or collecting personal information per the CCPA via its libraries as it advises app operators integrating those that they may “share with Meta Personal Information about California residents which would otherwise constitute a Sale or Sharing […].” In another example, per InMobi’s privacy policy (InMobi, 2024), “InMobi may be deemed to be ‘selling’ personal information, […]” indicating InMobi’s further downstream sale of California residents’ personal information after buying or collecting such from the app operators who integrate its libraries. We performed our evaluation conservatively. If a policy stated that a third party does not share or sell personal information per the CCPA, we accepted such statement as true. Our methodology could be applied to cover cases where the operator claims not to sell or share personal information but does so in practice, for example, as it appears from the operator’s app network traffic, possibly in combination with a privacy flag not set to opt-out. In such case the operator needs to either adapt the claim or the practice.
@@
To evaluate the likelihood to which an app in our app dataset is part of a business, we selected a random sample of 100 apps from our app dataset for determining whether their operators self-declared to run a “business” in their privacy policies. We performed this analysis on the 83 sample apps in our app dataset that had a Google Play Data Safety page with a link to their privacy policy and were still available at the time of the sampling. If an app’s Data Safety page linked to a privacy policy, we assumed that the linked privacy policy applies to the app. The privacy policy analysis for our GPC opt-out evaluation was also performed by the author with privacy law expertise.
@@
The setup and procedure for our GPC opt-out evaluation is similar to our UI opt-out evaluation, though, automated. After downloading the apps, we ran an automatic dynamic analysis for each app individually proceeding in batches for a total of two weeks with the analysis of one app taking about 5 minutes. We performed this analysis for the apps in our app dataset in October 2023. As shown in Figure 3, we instrumented the test device with a shell script, the controller, which ran on our laptop. The controller fetched an app from the app database on the laptop, installed it on the test device, ran it in the foreground under each of the four test conditions, and uninstalled it. If an app update was triggered right after installing an app, it would not have been possible to run the app immediately afterwards, which we check for.  [^21] Appendix 10.1 contains further details of our dynamic analysis with Appendix Figure 10 showing the procedure in detail.
@@
We want to find out if California residents can exercise their opt-out rights by sending GPC signals. To which extent do Android apps respect GPC signals despite the lack of platform-level API support for such? We are also interested in the effect of disabling the AdID. Thus, we tested each app under four conditions encompassing all possible combinations of access to the AdID (yes/no) and sending GPC signals (yes/no). Similar to previous work (Kollnig et al., 2021), for each condition, we captured each app’s network traffic for 30 seconds. We also added a timeout of 30 seconds before running an app’s first condition so that the app is fully initialized before capturing of the network traffic begins (Nguyen et al., 2022). The four test conditions under which we evaluated each app in our app dataset are as follows:
@@
    AdID: The app has access to the AdID. This condition is the base condition under which we expect apps to sell and share California residents’ personal information.
@@
    No AdID: The app does not have access to the AdID. Google prohibits targeting users with personalized advertising (Google, 2024c). If the app recognizes the AdID setting as a valid opt-out preference per the CCPA, it should notify ad networks of the opt-out via respective privacy flags.
@@
In the AdID + GPC and No AdID + GPC conditions we injected GPC headers into all requests, both HTTP and HTTPS requests, using Mitmproxy. In the AdID and No AdID conditions, we injected dummy headers into all requests to minimize any differences that might occur due to header injection in the other two conditions. In all four conditions the controller granted all permissions requested by an analyzed app at install time. [^22] Different from the UI Opt-out evaluation (Section 4.3), for the GPC opt-out evaluation we opted to grant apps location permission as we also wanted to explore location tracking. Appendix Figure 13 shows the top 50 permissions that the 1,811 apps in our app dataset declared in their manifest file. We tested whether granting or denying permissions made a difference for apps’ ad tracking behavior, which was not the case. When switching conditions, our controller cleared the persistent storage and SharedPreferences of the test device.
@@ 4.5. Limitations
Our study has various limitations. First, our evaluation of the applicability of the CCPA opt-out right, is based on privacy policies that app operators link from their app’s Play Store Data Safety page. It may be that a policy is not applicable to an app or a policy does not distinguish between multiple apps or the operator’s website, each of which may have different data practices. Further, by its very nature, dynamic analysis only runs code partially leaving us with an incomplete picture of apps’ behavior. Also, for our GPC opt-out evaluation we evaluated apps without user interaction, except for automatically enabling and disabling the AdID setting. [^23] Since automatically enabling or disabling the AdID may be impacted by pop-up screens or other interruptions, we manually re-analyzed apps that had unexpected AdID access and confirmed the results (Section 6.2.3). For determining opt-out effects we evaluate privacy flags, domain connections, and sharing of device identifiers. However, those are only proxies. For example, an app operator may simply disregard a user’s preference despite a privacy flag being set to opt-out. Our methodology does not allows us to observe server-to-server data sharing, for example, of device identifiers. Further, some apps we evaluated likely use encryption with certificate pinning. While we followed state-of-the-art methods for certificate-unpinning (Xiao et al., 2023; Kollnig et al., 2022a; Nguyen et al., 2022), we were likely not able to decrypt all encrypted traffic, which is an inherent limitation of any dynamic analysis of third-party code. For our identification of ad tracking domains we do not perform an independent evaluation and rely on previous research that led to creation of the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). Similar to previous work (Kumar et al., 2022), we use a commercial VPN, Mullvad, to simulate a California location so that the CCPA is applicable. Apps may identify our real location, e.g., via Mullvad’s IP address or our time zone, and behave differently. Thus, our results may not be an accurate reflection of app behavior for California residents. However, we are not aware of such behavior. Notably, in our GPC opt-out evaluation we did not notice different behavior in apps with and without GPS access.
@@ 5. UI Opt-out Evaluation
Main Result: Only 48 of the 100 apps we evaluated implement the legally mandated CCPA opt-out setting, suggesting non-compliance with the law.
@@ 5.2. Opt-out Effects
Generic/SDK/Cookie Opt-Out Setting 8 2.7%–13.3%
@@
Table 1. Counts of opt-out settings for the 100 apps for which we evaluated their opt-out UI. Figure 4 shows example UIs.
As Table 1 shows, only 48 out of 100 apps implement the legally mandated CCPA opt-out setting. The remaining 52 apps either have just a privacy policy link (45) or no opt-out setting or privacy policy link (7). As an example, calculating the confidence interval for the proportion of 45/100 (45%) apps with just a privacy policy link, using z-scores, indicates that with 95% confidence the true proportion of apps in our app dataset with only such link is 35%–55%. Thus, even without evaluating any network traffic, our results suggest the existence of non-compliance. While apps in some game categories tend to have higher percentages of non-compliant apps, non-compliance is generally observable across all app categories (Appendix Table 3).
Turning to our network traffic analysis, some apps did not reflect the opted out state in the privacy flags they passed to third parties. Thus, even if California residents opt out, those apps are still selling and sharing personal information. Out of the 14 apps that included the third-party ad network Vungle (Vungle, 2025), two apps still set the ccpa status flag to opted_in in the opted out condition so that Vungle continued to collect and buy personal information. Similarly, one app also still had Google’s Restricted Data Processing (RDP) (Google, 2025) set to 0. These results show that opt-outs are not always respected as app operators do not notify third parties of the opt-out so that they are in actuality not opted out. We did not see differences for connecting to ad tracking domains and disclosing device identifiers between opted out and not opted out conditions.
@@ 6. GPC Opt-out Evaluation
Main Result: While we estimate with 95% confidence that 62%–81% of apps in our app dataset must respect the CCPA’s opt-out right, many apps do not do so. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out.
@@ 6.1. CCPA Opt-out Right Applicability
Evaluating how many apps are selling or sharing personal information, we find that 1,787/1,811 (99%) of the apps in our app dataset connect to at least one of the most prevalent 29 ad tracking domains, each of which has a privacy policy stating to (1) buy or collect or (2) sell or share personal information per the CCPA (Figure 5). [^24] Performing the evaluation for all third-party tracking domains on the X-Ray Tracker List only increases the coverage by two apps to 1,789/1,811 (99%), which is indicative for the industry concentration in the online ad industry. Further, for the random sample of 83 apps (Section 4.4), 60/83 (72%) of their policies implied that the app operators run a business per the CCPA by mentioning the applicability of the CCPA or the explicit selling or sharing of personal information (Figure 6). Thus, we estimate that 17870.72=1,2871787*0.72=1,287 apps in our app dataset are subject to the CCPA opt-out right. Calculating the confidence interval for the proportion of 60/83 (72%) sample apps, using z-scores, indicates that with 95% confidence the true proportion of app operators running a business is 63%–82%. Applying this range to the 1,787 apps that connect to at least one of the 29 most prevalent tracking domains, we arrive at an estimate of 1,126 to 1,465 (62%–81%) apps in our app dataset being subject to the CCPA opt-out right.
@@ 6.2. Opt-out Effects
We now evaluate the effectiveness of sending GPC signals for opting out on Android apps per the CCPA and what effect, if any, disabling the AdID has. Overall, we find little evidence that the apps in our app dataset respect GPC. Notably, we see no substantial impact of GPC on apps’ notification to third parties about our opt-out via the IAB’s US Privacy String (IAB, 2024) and company-specific privacy flags (Section 6.2.1). We also do not observe a decrease in the number of connections to ad tracking domains (Section 6.2.2) or the disclosure of device identifiers (Section 6.2.3).
6.2.1. Sending CCPA Privacy Flags to Ad Networks
The IAB provides an industry-wide CCPA privacy flag in the form of the US Privacy String (IAB, 2024), to be implemented as an HTTP cookie or a JavaScript property. [^25] On January 31, 2024 the US Privacy String was deprecated and integrated into the IAB’s Global Privacy Platform (GPP) that covers privacy flags for multiple US states and international jurisdictions (IAB Tech Lab, 2022a). The IAB now recommends use of GPP instead of the US Privacy String. GPP contains the usca California privacy section corresponding to the US Privacy String (IAB Tech Lab, 2022b). Since we performed our analysis before the deprecation date we focus here on the US Privacy String. Both the US Privacy String as well as its GPP equivalent can be stored on Android devices using SharedPreferences for the app context (IAB Tech Lab, 2020, 2022a). We find that when opting out via GPC, privacy flags generally do not change to reflect the CCPA opt-out status. However, some apps opted us out regardless of whether or not we sent GPC signals or enabled or disabled the AdID.
@@ 6.2.1. Sending CCPA Privacy Flags to Ad Networks
For 199/1,811 (11%) apps in our app dataset we found a US Privacy String key, us_privacy, and a corresponding US Privacy String value in their network traffic. However, the value of the US Privacy String was rarely set to opt-out, that is, rarely had a Y in its third position, which would indicate a “Yes” for a California resident’s status of being opted out (IAB, 2024) (Figure 7). Disabling the AdID did not lead to substantially different opt-out rates as the comparison between the AdID and No AdID conditions shows (Appendix Figure 14). Overall, the rate of 1YY- US Privacy String values, which indicate that a California resident’s opt-out right is respected, is low regardless of the tested condition. Our results suggest that turning on GPC does not help California residents to exercise their CCPA opt-out right as measured via the US Privacy String. Disabling the AdID did not change this result.
@@
Table 2. Counts and percentages of apps that share privacy flags with ad networks. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out, which corresponds to only 7% of the total apps.
@@
Our evaluation of company-specific privacy flags further confirms our findings. Table 2 shows counts and percentages of apps in our app dataset with left-column values indicating company-specific privacy flag opt-in status, middle-column values indicating opt-out status, and right-column values indicating opt-out percentages. Just as for the US Privacy String, company-specific privacy flags rarely changed when disabling the AdID or sending GPC signals. Generally, the apps for which we detected privacy flags with opt-out values had this status regardless of the GPC and AdID setting. It seems that apps opt out all California residents by default or trigger the opt-out status when they detect requests from a California IP address. It is noteworthy that Google’s rdp flag was more often set to opt-out than to opt-in, albeit, at comparatively low absolute levels. Generally it is evident that sending GPC signals does not make a substantial difference overall. Sending GPC signals does not have a major effect for opting out California residents per the CCPA as measured via company-specific privacy flags. Disabling the AdID does not lead to a different result either.
@@
There are 43 apps in our app dataset that implement OneTrust (OneTrust, 2025), a popular CMP, we analyzed under which conditions these apps set OneTrust’s IsGpcEnabled flag to True. Apps’ HTTP responses for setting the flag fall into different categories. However, since they are not using consistent category names, the categorization of the IsGpcEnabled flag is not directly possible. Therefore, we identified and manually coded the observed category names into five overarching opt-out categories, similar to cookie categories: Functional, Performance, Social Media, Strictly Necessary, and Targeted Advertising. After our categorization, consistent with our previous results, we observed no substantial differences in the IsGpcEnabled flag values across the four conditions (Figure 8). [^26] Most apps received a single response from OneTrust except com.greencopper.countryjam, which returned two identical responses that we, thus, only included once in the final count. One app, com.disney.wdw.android, did not provide a OneTrust response for the AdID + GPC condition, indicated by the shorter red bars in Figure 8. A few apps came with additional category names that did not fit any of our categories. We omitted those from the final count for the respective apps.

### v6

@@ 4.4. GPC Opt-out Evaluation
In addition to our UI opt-out evaluation, we performed a GPC opt-out evaluation. The “selling” or “sharing” of personal information by an app operator corresponds to a third party’s “buying” or “collecting” of personal information. Thus, to determine whether an app is selling or sharing personal information we identified the most prevalent third-party tracking domains across all apps in our app dataset. Our starting point for this task was the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). We found that a set of 29 third-party tracking domains from the X-Ray Tracker List lead to a 99% coverage with at least one tracking domain—whose privacy policy allowed to (1) buy or collect or (2) sell or share personal information per the CCPA—in each app in our app dataset (Section 6.1). [^20] We combined trackers with different subdomains in our app dataset into one domain to match the tracker domains in the X-Ray Tracker List. For example, we combined app.adjust.com, view.adjust.com, app.us.adjust.com, and s2s.adjust.com into adjust.com. For the privacy policy evaluation we included developer documentation and looked for explicit references that—per the CCPA—allows the buying or collecting of personal information from app operators or the selling or sharing of such.
@@
To illustrate, Meta’s policy (Meta, 2025) indicates that it is buying or collecting personal information per the CCPA via its libraries as it advises app operators integrating those that they may “share with Meta Personal Information about California residents which would otherwise constitute a Sale or Sharing […].” In another example, per InMobi’s privacy policy (InMobi, 2024), “InMobi may be deemed to be ‘selling’ personal information, […]” indicating InMobi’s further downstream sale of California residents’ personal information after buying or collecting such from the app operators who integrate its libraries. We performed our evaluation conservatively. If a policy stated that a third party does not share or sell personal information per the CCPA, we accepted such statement as true. Our methodology could be applied to cover cases where the operator claims not to sell or share personal information but does so in practice, for example, as it appears from the operator’s app network traffic, possibly in combination with a privacy flag not set to opt-out. In such case the operator needs to either adapt the claim or cease the practice.
@@
To evaluate the likelihood to which an app in our app dataset is part of a business, we selected a random sample of 100 apps from our app dataset for determining whether their operators self-declared to run a “business” in their privacy policies. We performed this analysis on the 83 sample apps in our app dataset that had a Google Play Data Safety page with a link to their privacy policy and were still available at the time of the sampling. Given that an app’s Data Safety page linked to a privacy policy, we assumed that the linked policy applies to the app.
@@
The setup and procedure for our GPC opt-out evaluation is similar to our UI opt-out evaluation, though, automated. After downloading the apps, we ran an automated dynamic analysis for each app individually proceeding in batches for a total of two weeks with the analysis of one app taking about 5 minutes. We performed this analysis for the apps in our app dataset in October 2023. As shown in Figure 3, we instrumented the test device with a shell script, the controller, which ran on our laptop. The controller fetched an app from the app database on the laptop, installed it on the test device, ran it in the foreground under four test conditions, and uninstalled it. If an app update was triggered right after installing an app, it would not have been possible to run the app immediately afterwards, which we checked for.  [^21] Appendix 11.1 contains further details of our dynamic analysis with Appendix Figure 10 showing the procedure in detail.
@@
We wanted to find out if California residents can exercise their opt-out rights by sending GPC signals. To which extent do Android apps respect GPC signals despite the lack of platform-level API support for such? We were also interested in the effect of disabling the AdID. Thus, we tested each app under four conditions encompassing all possible combinations of access to the AdID (yes/no) and sending GPC signals (yes/no). Similar to previous work (Kollnig et al., 2021), for each condition, we captured each app’s network traffic for 30 seconds. We also added a timeout of 30 seconds before running an app’s first condition so that the app is fully initialized before capturing of the network traffic begins (Nguyen et al., 2022). The four test conditions under which we evaluated each app in our app dataset are the following:
@@
    AdID: The app has access to the AdID. This condition is the base condition under which we expect some apps to sell and share California residents’ personal information.
@@
    No AdID: The app does not have access to the AdID. Google prohibits targeting users with personalized advertising (Google, 2025e). If the app recognizes the AdID setting as a valid opt-out preference per the CCPA, it should notify integrated ad networks of the opt-out via respective privacy flags.
@@
In the AdID + GPC and No AdID + GPC conditions we injected GPC headers into all requests, for both HTTP and HTTPS, using Mitmproxy. In the AdID and No AdID conditions, we injected dummy headers into all requests to minimize any differences that might occur due to header injection in the other two conditions. In all four conditions the controller granted all permissions requested by an analyzed app at install time. [^22] Different from the UI Opt-out evaluation (Section 4.3), for the GPC opt-out evaluation we opted to grant apps location permission as we also wanted to explore location tracking. Appendix Figure 13 shows the top 50 permissions that the 1,811 apps in our app dataset declared in their manifest file. We tested whether granting or denying permissions made a difference for apps’ ad tracking behavior, which was not the case. When switching conditions, our controller cleared the persistent storage and SharedPreferences of the test device.
@@ 4.5. Limitations
Our study has various limitations. First, our evaluation of the applicability of the CCPA opt-out right is based on privacy policies that app operators link from their app’s Play store Data Safety page. It may be that a policy is not applicable to an app or a policy does not distinguish between multiple apps or the operator’s website, each of which may have different data practices. Further, by its very nature, dynamic analysis only runs code partially leaving us with an incomplete picture of apps’ behavior. Also, for our GPC opt-out evaluation we evaluated apps without user interaction save for automatically enabling and disabling the AdID setting. [^23] Automatically enabling or disabling the AdID may be impacted by pop-up screens or other interruptions. We encountered some apps with unexpected AdID access. We manually re-analyzed these apps and confirmed the results (Section 6.2.3). For determining opt-out effects we evaluated privacy flags, domain connections, and sharing of device identifiers. However, those are only proxies. For example, an ad network may simply disregard a user’s right despite a privacy flag being set to opt-out. Our methodology does not allow us to observe server-to-server data sharing, for example, of device identifiers. Further, some apps we evaluated likely used encryption with certificate pinning. While we followed state-of-the-art methods for certificate-unpinning (Xiao et al., 2023; Kollnig et al., 2022a; Nguyen et al., 2022), we were likely not able to decrypt all encrypted traffic, which is an inherent limitation of any dynamic analysis of third-party code. For our identification of ad tracking domains we did not perform an independent evaluation and relied on previous research that led to creation of the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). Similar to previous work (Kumar et al., 2022), we used a commercial VPN, Mullvad, to simulate a California location so that the CCPA is applicable. Apps may have identified our real location, e.g., via Mullvad’s IP address or our time zone setting, and behaved differently. Thus, our results may not be an accurate reflection of app behavior for California residents. However, we are not aware of such behavior. Notably, in our GPC opt-out evaluation we did not notice different behavior in apps with and without GPS access.
@@ 5. UI Opt-out Evaluation
Main Result: Only 48 of the 100 apps we evaluated implement the legally mandated CCPA opt-out setting, which suggests that CCPA opt-out right compliance on the Android platform is generally low.
@@ 5.2. Opt-out Effects
Generic/SDK/Cookie Opt-Out Setting 8 3%–13%
@@
Table 1. Counts of opt-out settings for the 100 apps for which we performed the UI opt-out evaluation and confidence intervals for the prevalence of respective settings in our app dataset. Figure 4 shows example UIs.
As Table 1 shows, only 48 out of 100 apps implement the legally mandated CCPA opt-out setting. The remaining 52 apps either have just a privacy policy link (45) or no opt-out setting or privacy policy link (7). As an example, calculating the confidence interval for the proportion of 45/100 (45%) apps with just a privacy policy link, using z-scores, indicates that with 95% confidence the true proportion of apps in our app dataset with only such link is 35%–55%. Thus, even without evaluating any network traffic, our results suggest the broad existence of non-compliance. While apps in some game categories tend to have higher percentages of non-compliant apps, non-compliance is generally observable across all app categories (Appendix Table 3).
Turning to our network traffic analysis, some apps did not reflect the opted out state in their privacy flags. Thus, even if California residents opt out, those apps are still selling and sharing personal information. Out of the 14 apps that included the third-party ad network Vungle (Vungle, 2025), two apps still set the ccpa status flag to opted_in in the opted out condition so that Vungle continued to collect and buy personal information. Similarly, one app also still had Google’s Restricted Data Processing (RDP) (Google, 2025b) set to 0. These results show that opt-out rights are not always respected as app operators do not notify third parties of users’ opt-outs so that those remain opted in. We did not see differences for connecting to ad tracking domains and disclosing device identifiers between opted out and not opted out conditions.
@@ 6. GPC Opt-out Evaluation
Main Result: While we estimate with 95% confidence that 62%–81% of apps in our app dataset must respect the CCPA opt-out right, many apps do not do so. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted_in while only 26 had set it to opted_out.
@@ 6.1. CCPA Opt-out Right Applicability
Evaluating how many apps are selling or sharing personal information, we find that 1,787/1,811 (99%) of the apps in our app dataset connect to at least one of the most prevalent 29 ad tracking domains with a privacy policy stating to (1) buy or collect or (2) sell or share personal information per the CCPA (Figure 5). [^24] Performing the evaluation for all third-party tracking domains on the X-Ray Tracker List only increases the coverage by two apps to 1,789/1,811 (99%), which highlights the concentration of companies in the online ad industry. Further, for the random sample of 83 apps (Section 4.4), 60/83 (72%) of their policies implied that the app operators run a business per the CCPA by mentioning the applicability of the CCPA or the explicit selling or sharing of personal information (Figure 6). Thus, we estimate that 1,7870.72=1,2871,787*0.72=1,287 apps in our app dataset are subject to the CCPA opt-out right. Calculating the confidence interval for the proportion of 60/83 (72%) sample apps, using z-scores, indicates that with 95% confidence the true proportion of app operators running a business is 63%–82%. Applying this range to the 1,787 apps that connect to at least one of the 29 most prevalent tracking domains, we arrive at an estimate of 1,126 to 1,465 (62%–81%) apps in our app dataset being subject to the CCPA opt-out right.
@@ 6.2. Opt-out Effects
Overall, we find little evidence that the apps in our app dataset respect GPC. Notably, we observe no substantial impact of GPC on apps’ notifications to third parties via the IAB’s US Privacy String (IAB Tech Lab, 2025) and company-specific privacy flags (Section 6.2.1). We also do not observe a decrease in the number of connections to ad tracking domains (Section 6.2.2) or the disclosure of device identifiers (Section 6.2.3).
6.2.1. Setting CCPA Privacy Flags
The IAB provides an industry-wide CCPA privacy flag in the form of the US Privacy String (IAB Tech Lab, 2025), to be implemented as an HTTP cookie or a JavaScript property. [^25] On January 31, 2024 the US Privacy String was deprecated and integrated into the IAB’s Global Privacy Platform (GPP) that covers privacy flags for multiple US states and international jurisdictions (IAB Tech Lab, 2022). The IAB now recommends use of GPP instead of the US Privacy String. GPP contains the usca California privacy section corresponding to the US Privacy String (IAB Tech Lab, 2024a). Since we performed our analysis before the deprecation date we focus here on the US Privacy String. Both the US Privacy String as well as its GPP equivalent can be stored on Android devices using SharedPreferences for the app context (IAB Tech Lab, 2024b, 2022). We find that when opting out via GPC, privacy flags generally do not change to reflect the CCPA opt-out status. However, some apps opt out users regardless of whether they receive GPC signals or have access to the user’s AdID.
@@ 6.2.1. Setting CCPA Privacy Flags
For 199/1,811 (11%) apps in our app dataset we found a US Privacy String key, us_privacy, and a corresponding US Privacy String value in their network traffic. However, the value of the US Privacy String was rarely set to opt-out, that is, rarely had a Y in its third position, which would indicate a “Yes” for a California resident’s status of being opted out (IAB Tech Lab, 2025) (Figure 7). Also, disabling the AdID did not lead to substantially different opt-out rates as the comparison between the AdID and No AdID conditions shows (Appendix Figure 14). Overall, the rate of 1YY- US Privacy String values, which indicate that a California resident’s opt-out right is respected, is low regardless of the tested condition. Our results suggest that turning on GPC does not help California residents to exercise their CCPA opt-out right as measured via the US Privacy String. Disabling the AdID did not change this result.
Google Supersonic/Unity Vungle
@@
Table 2. Counts and percentages of apps that set company-specific privacy flags. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted_in while only 26 had set it to opted_out, which corresponds to only 26/364 (7%) of the total apps.
@@
Our evaluation of company-specific privacy flags further confirms our findings. We evaluated privacy flags of three companies: Google, Supersonic/Unity, and Vungle. Table 2 shows counts and percentages of apps in our app dataset with left-column values indicating company-specific privacy flags set to opt-in status, middle-column values indicating opt-out status, and right-column values indicating opt-out percentages. Just as for the US Privacy String, company-specific privacy flags rarely changed when disabling the AdID or sending GPC signals. Generally, the apps for which we detected privacy flags with opt-out values had this status regardless of the GPC and AdID setting. It seems that apps opt out all California residents by default or trigger the opt-out status when they detect requests from a California IP address. Notably, Google’s rdp flag was more often set to opt-out than to opt-in, albeit, at comparatively low levels. Generally it is evident that sending GPC signals does not make a substantial difference overall. Sending GPC signals does not have a major effect for opting out California residents per the CCPA as measured via company-specific privacy flags. Disabling the AdID does not lead to a different result either.
@@
There are 43 apps in our app dataset that implement OneTrust (OneTrust, 2025), a popular CMP. For each app we analyzed under which condition it sets OneTrust’s IsGpcEnabled flag to True. Apps’ HTTP responses for setting the flag fall into different categories. However, since app operators do not use consistent category names, the categorization of the IsGpcEnabled flag is not directly possible. Therefore, we identified and manually coded the observed category names into five overarching opt-out categories, similar to cookie categories: Functional, Performance, Social Media, Strictly Necessary, and Targeted Advertising. After our categorization, consistent with our previous results, we observed no substantial differences in the IsGpcEnabled flag values within each category across the four conditions (Figure 8). [^26] Most apps received a single response from OneTrust except com.greencopper.countryjam. This app received two identical responses, which we counted as one response. One app, com.disney.wdw.android, did not receive a OneTrust response for the AdID + GPC condition, indicated by the shorter red bars in Figure 8. A few apps’ categories did not fit any of our categories. We omitted those from our analysis. Notably, the IsGpcEnabled flag values for Targeted Advertising did not change substantially across conditions.